OpenStudy (anonymous):

write 6 (cos 330 deg +i sin 330deg in rectagular form

OpenStudy (anonymous):

Use eulers formula, and what you know about trigonometric functions

OpenStudy (anonymous):

First put it into exponential form

OpenStudy (anonymous):

\[re^{i\phi}=r(\cos(\phi)+i\sin(\phi))\]

OpenStudy (anonymous):

I assume you want to plot this on an argand diagram? where the imaginary line corresponds to your y axis and the real line corresponds to your horizontal axis

OpenStudy (anonymous):

no they dnt want any diagram

OpenStudy (anonymous):

Then how do you expect to graph a complex number in the real Cartesian plane

OpenStudy (anonymous):

is it -3 sqrt 3+3i

OpenStudy (anonymous):

Oh you just want to simplify it in terms of radicals

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I would first change your degrees to radians, and then calculate the cosine and sine functions using memorized values.

OpenStudy (anonymous):

alternatively you could just type it on this site, www.wolframalpha.com

OpenStudy (anonymous):

It should give you a list of expressions all equivilent

OpenStudy (anonymous):

Here il do it for you, and show you what I get

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ye

OpenStudy (mww):

6 (cos 330 deg +i sin 330deg) Rectangular means a+ ib form...but hey! It does look a bit like it already, so simplify that! Expand: 6cos330 + 6i sin330 Now cos 330 = cos 30 = sqrt(3)/2(using ASTC rules) and sin330 = - sin30 = -1/2 So the answer is 6(sqrt(3)/2) + 6i (-1/2) = \[6\sqrt{3}/2 - 3i = 3\sqrt{3} - 3i\]

OpenStudy (anonymous):

ohhh i c thanks